The credit remaining on a phone card (in dollars) is a linear function of the total calling time made with the card (in minutes). The remaining credit after
25
minutes of calls is
$26.50
, and the remaining credit after
53
minutes of calls is
$22.58
. What is the remaining credit after
61 minutes of calls ?

Answer by mathmate(429)

Question:
The credit remaining on a phone card (in dollars) is a linear function of the total calling time made with the card (in minutes). The remaining credit after
25 minutes of calls is $26.50, and the remaining credit after 53 minutes of calls is $22.58. What is the remaining credit after 61 minutes of calls ?
Solution:
Given linear function, we have two given points of
f(n), namely
(25,26.50), (53,22.58)
So apply the two point formula:
slope m=(y1-y1)/(x2-x1)
=(22.58-26.5)/(53-25)
=-3.92/28
=-0.14

the equation required is
(y-y1)=m(x-x1)
y-26.5=-.14(x-25)
simplify to get
y=-0.14x+30

The amount left after 61 calls
A=f(61)=-0.14*61+30=21.46